# Practitioners’ Discussions on Building LLM-based Applications for Production

### v1

@@ IV-B Co-Occurrences of Themes and Topics in Videos
We saw some intersections among topics that fall into the same architecture style. Within a RAG system, prompt engineering is a crucial part (17 joint videos). It specifies how to incorporate the additional context coming from the retrieval part, prepares the LLM for factual answers, and may even rewrite the user prompt to a more suitable phrasing. However, as discussed previously, it is crucial to be able to evaluate the effects of incorporating RAG components. Hence, we observe also that evaluation is often mentioned together, with in total 21 joint videos — the highest number of co-occurring topics. With similar importance of evaluation for Fine-Tuning (a model; 18) and Prompt Engineering (17).
Some notable further aspects that come with the development of a RAG system are Application Security (11), Context Length (9), Latency (10), Memory (9), Model Size and Performance (11), On-premise or Cloud (11), and User Interfaces (11). So, by just looking at the corresponding row, practitioners can quickly see what further topics may play a role when considering to implement a RAG system.
@@
From our analysis on co-occurrences of topics, we can infer, at least, three interesting observations: topics to consider when choosing an architecture, alternatives of techniques, and tools for realization. Building upon such a matrix has the potential to become a valuable tool for practitioners in the future for quickly obtaining an overview of the field and find relevant content. Opening this way is a notable contribution to the field.
@@ V Discussion
The breadth of the topics is huge. We observe that practitioners exchange architecture styles of LLM-based applications, discuss technical challenges as for fine-tuning and prompt engineering, and highlight very practical aspects, such as risks, ethics, security, and costs. At least for software engineering venues, we do not see such a rich set of topics. Our overview can clearly pinpoint some lack in practice. Notably, the ability to systematically evaluate prompts and RAG systems seem to be one of the biggest concerns, even despite intensive research on that topic [35, 36, 27]. It is unclear whether academic approaches are unknown to industry or whether we miss some practical aspects in our research to make an impact. This is an interesting avenue for the future.
@@ V-A Threats to Validity
A threat to external validity or generalization we come from the selection process of the videos to obtain our initial set. Being aware of this possible threat, we applied two search strategies: First, we search for playlists coming from high professional communities (e.g., MLOps.community) to include industry conference talks about that topic and obtain high quality, industry-relevant material; second, we used keywords to select videos thematically relevant, but without such a community bias. Given the breath of the topics, we mitigated the threat to external validity.
@@ Acknowledgment
The authors acknowledge the financial support by the Federal Ministry of Education and Research of Germany and by Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus in the programme Center of Excellence for AI-research “Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig”, project identification number: ScaDS.AI. Siegmund’s work has been funded by the German Research Foundation (SI 2171/2-2).

### v2

@@ IV-B Co-Occurrences of Themes and Topics in Videos
We saw some intersections among topics that fall into the same architecture style. Within a RAG system, prompt engineering is a crucial part (17 joint videos). It specifies how to incorporate the additional context coming from the retrieval part, prepares the LLM for factual answers, and may even rewrite the user prompt to a more suitable phrasing. However, as discussed previously, it is crucial to be able to evaluate the effects of incorporating RAG components. Hence, we observe also that evaluation is often mentioned together with RAG-systems, with in total 21 joint videos—the highest number of co-occurring topics. With similar importance of evaluation for Fine-Tuning (a model; 18) and Prompt Engineering (17).
Some notable further aspects that come with the development of a RAG system are Application Security (11), Context Length (9), Latency (10), Memory (9), Model Size and Performance (11), On-premise or Cloud (11), and User Interfaces (11). So, by just looking at the corresponding row in Figure 3, practitioners can quickly see what further topics may play a role when considering to implement a RAG system.
@@
From our analysis on co-occurrences of topics, we can infer, at least, three interesting observations: (1) topics to consider when choosing an architecture, (2) alternatives of techniques, and (3) tools for realization. Building upon such a matrix has the potential to become a valuable tool for practitioners in the future for quickly obtaining an overview of the field and find relevant content. Introducing this initial matrix is a notable contribution to the field.
@@ V Discussion
The breadth of the topics is huge. We observe that practitioners exchange architecture styles of LLM-based applications, discuss technical challenges as for fine-tuning and prompt engineering, and highlight very practical aspects, such as risks, ethics, security, and costs. At least for software engineering venues, we do not see such a rich set of topics. Our overview can clearly pinpoint different relevant topics in practice. Notably, the ability to systematically evaluate prompts and RAG systems seem to be one of the biggest concerns, even despite intensive research on that topic [35, 36, 27]. It is unclear whether academic approaches are unknown to industry or whether we miss some practical aspects in our research to make an impact. This is an interesting avenue for the future.
@@ V-A Threats to Validity
A threat to external validity or generalization comes from the selection process of the videos to obtain our initial set. Being aware of this possible threat, we applied two search strategies: First, we search for playlists coming from professional communities (e.g., MLOps.community) and industry conference talks about that topic and obtain high quality, industry-relevant material; second, we used keywords to select videos thematically relevant, but without such a community bias. Given the breadth of the topics, we mitigated the threat to external validity.
@@ Acknowledgment
The authors acknowledge the financial support by the Federal Ministry of Education and Research of Germany and by Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus in the programme Center of Excellence for AI-research “Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig”, project identification number: ScaDS.AI. Siegmund’s work has been funded by the German Research Foundation (SI 2171/2-2). This work is supported by the European Social Fund (ESF) and the German state of Saxony under grant number A.100760692 (Project: Teaching-AI).